OpenStudy (anonymous):

Find the zeros of 3x^2-12x+2=0 using the quadratic formula

OpenStudy (anonymous):

the solutions are 3.825 and 0.174

OpenStudy (anonymous):

How did you get that??

OpenStudy (anonymous):

using the quadratic formulae

OpenStudy (anonymous):

No frick -_-. I need to know the steps. I don't want a direct answer

OpenStudy (anonymous):

if the quadratic equation is of the form ax^2 + bx + c = 0 then its roots(or in other words zeros ) are given by\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

I KNOW THAT. I'm wanting to understand the steps that you took in order to get 3.825 and 0.174. It seems very clear that you're trying to avoid helping me understand. @sourav_aich

OpenStudy (anonymous):

here a = 3 b = -12 c = 2 put these values in the equation you will get the answer use a calculator

OpenStudy (anonymous):

No. In fact I won't. If you weren't wanting to actually help in the first place- you shouldn't have. I don't freaking care if I can get the answer through a calculator. Intelligent people actually like to KNOW how to solve it

OpenStudy (anonymous):

@sourav_aich

OpenStudy (anonymous):

then you can also do without the use of calculator just put the values in the equation and calculate

OpenStudy (anonymous):

Why would I even post this on OS if I knew how to solve it genius -_-

OpenStudy (anonymous):

after putting the values in the formula you will get an expression like this \[\frac{ 12\pm \sqrt{12^{2}-4*3*2} }{ 2*3 }\] \[\frac{ 12 \pm \sqrt{120}}{ 6 }\] \[\frac{ 12 \pm 2\sqrt{30}}{ 6 }\] \[\frac{ 6\pm \sqrt{30} }{ 3 }\] so the solutions are \[\frac{ 6+\sqrt{30} }{ 3 } and \frac{ 6+\sqrt{30} }{ 3 } \]

OpenStudy (anonymous):

That's all I was asking for! Awesome